OpenStudy (anonymous):

what property is 1x+1?

OpenStudy (amistre64):

dunno, how many properties we got?

OpenStudy (anonymous):

identity property of multiplication

OpenStudy (anonymous):

1 times something is itself.

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

deff

OpenStudy (anonymous):

either that or the distributive property lol

OpenStudy (anonymous):

i need to know for sure

OpenStudy (anonymous):

both would work, you just need to know how to justify it.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

^.^

OpenStudy (anonymous):

what property is x+1?

OpenStudy (amistre64):

.... the add by 1 property? i dont think there is one

OpenStudy (anonymous):

i think its commutative

OpenStudy (amistre64):

not quite; comm would be x=1 = 1+x

OpenStudy (amistre64):

but without the typo

OpenStudy (amistre64):

x+1 all alone is no special property other then simply adding

OpenStudy (anonymous):

1(x+1) = x+1 Identity property.. or distributive.

OpenStudy (anonymous):

k anyways...t hanks for ur help

OpenStudy (anonymous):

kk